# Getting to the Bottom of Serverless Billing

### v1

@@ Abstract.
Public cloud serverless platforms have attracted a large user base due to their high scalability, plug-and-play deployment model, and pay-per-use billing. However, compared to virtual machines and container hosting services, modern serverless offerings typically impose higher per-unit time and resource charges. Additionally, billing practices such as wall-clock allocation-based billing, invocation fees, and usage rounding up can further increase costs.
This work, for the first time, holistically demystifies these costs by conducting an in-depth, top-down characterization and analysis from user-facing billing models, through request serving architectures, and down to operating system scheduling on major public serverless platforms. We quantify, for the first time, how current billing practices inflate billable resources up to 5.49×5.49$\times$5.49 × beyond actual consumption. Also, our analysis reveals previously unreported cost drivers, such as operational patterns of serving architectures that create overheads, details of resource allocation during keep-alive periods, and OS scheduling granularity effects that directly impact both performance and billing. By tracing the sources of costs from billing models down to OS scheduling, we uncover the rationale behind today’s expensive serverless billing model and practices and provide insights for designing performant and cost-effective serverless systems.
footnotetext: * Ma is affiliated with Johns Hopkins University, but conducted the research at the University of British Columbia.
@@ 1. Introduction
Serverless computing has become one of the mainstream cloud computing paradigms, enabling developers to quickly deploy scalable and event-driven applications on the cloud without needing to manage the underlying infrastructure (Jonas et al., 2019; web, 2023b). Major cloud providers offer serverless computing solutions, such as AWS Lambda (web, 2025w), Google Cloud (GCP) Run functions (web, 2025i), Azure Functions (web, 2025e), IBM Cloud Code Engine (web, 2025p), and Cloudflare Workers (web, 2025j). Serverless computing stands out as the purest existing pay-per-use cloud model, offering automated scaling—from zero to thousands of instances in seconds—and fine-grained billing. As a result, it is often advertised as cost-efficient (Eismann et al., 2021b; web, 2025v; Jonas et al., 2019; Wood, 2024; Cloudflare, 2025a).
@@
The widely acknowledged benefits of serverless architectures—such as high scalability, fine-grained pay-per-use billing, freedom from infrastructure management, and seamless integration with other cloud services—are not without associated costs (Eivy and Weinman, 2017; Hellerstein et al., 2018; Wen et al., 2023). In terms of the per-unit resource price, serverless offerings are often priced higher than other cloud computing paradigms, such as virtual machines (VMs) and containers running on container hosting platforms. We demonstrate this by comparing the price of AWS Lambda, AWS EC2 VMs, and AWS Fargate containers, all configured on identical ARM-based hardware in the us-east-2 region. We specifically chose ARM due to the diverse and performance-varying nature of AWS’s x86 processors, which complicates fair comparisons across services. An AWS Lambda function with 1 vCPU, 1,769 MB of memory, and 512 MB of ephemeral storage costs \$$\$2.3034\times 10^{-5}$ per second (web, 2025v), while a compute-optimized EC2 instance (c6g.medium) with 1 vCPU, 2 GB memory, and 1 GB storage and an AWS Fargate container with the identical resource allocation as EC2 cost only \$$\$9.4753\times 10^{-6}$ and \$$\$1.1003\times 10^{-5}$ per second, which are 41.1% and 47.8% of the AWS Lambda price. The cost of VMs can be further decreased by at least two times if using a burstable instance (e.g., AWS EC2 t4g.small flavor). Also, this comparison does not include the invocation fee of AWS Lambda, which is \$$\$2\times 10^{-7}$ for each request, whereas EC2 instances and Fargate containers do not charge request fees. Additionally, our analysis of billing practices on major serverless platforms uncovers significant over-accounting (§2), showing that users can be charged for computing resources up to 5.49×5.49$\times$5.49 × greater than their actual usage.
@@
These observations motivate a fundamental research question: What makes serverless expensive? We argue that the root cause of the high unit prices and expensive billing practices in serverless lies in the architecture of modern serverless computing systems. Resource consumption and overhead incurred by the underlying runtimes and control plane for request serving, such as sandbox provisioning, isolation, request dispatch, and keep-alive (KA), translate into higher per-unit charges passed on to serverless users. Additionally, some of our measurements of resource allocation patterns and performance behaviors on major serverless platforms point to part of the execution costs and performance fluctuations to the underlying operating system (OS) scheduling mechanisms.
@@
Serverless Platform Billable Time Billable Resources1 Billing Granularity/Cutoffs Control Knobs and Steps
AWS Lambda (web, 2025w, k, v)
@@
(Request-Based Billing) (web, 2025i, t)
@@ 1.5 GB memory and 1 vCPU)
Huawei Cloud Function Graph (web, 2025o)
@@
1This table and related analysis in §2 focus on the most basic billable computing resources (i.e., CPU and memory). Other billable resources (e.g., storage, GPUs, and network bandwidths) may apply in practice. 2AWS bills wall-clock turnaround time that includes cold start latency (initialization duration) starting August 2025 (Gupta and Gebhart, 5 07).
Table 1. The notion of billable time, billable resources, and billing granularity varies across different serverless platforms (as of 2025-05-15).
To uncover these costs, a detailed analysis of current billing models together with measurements of the underlying serverless systems is required. Previous studies have characterized major serverless platforms in terms of architecture, performance, and resource management (Adzic and Chatley, 2017; Wang et al., 2018; Yu et al., 2020; Cvetković et al., 2023). However, as serverless computing evolves rapidly and more serverless offerings become available, some earlier measurements do not reflect or fully capture the latest billing scheme and operation patterns (e.g., serving architecture and KA behaviors) of the public serverless computing platforms. In this work, we revisit some of the previous measurements and extend some of their performance and overhead characterization to fit modern serverless systems.
We adopt a top-down approach to analyze serverless costs. We start with user-facing billing models and conduct large-scale trace analysis on the billing scheme. Then, we analyze the performance, overhead, and resource allocation patterns of modern serverless request serving architectures. Finally, we investigate the impact of OS scheduling in detail. By tracing sources of costs from the billing model down to kernel scheduling, we provide the first comprehensive decomposition of serverless overhead and reveal the rationale behind current billing practices. For example, we observe new billing models adopted by providers, investigate modern serverless serving architectures (§3.1 and §3.2), and track KA policy changes (§3.3). We also report multiple new aspects for the first time, such as large-scale, trace-based billing model analysis (§2), details of resource allocation during KA (§3.3), and the impact of OS-level scheduling granularity (§4). Specifically, our main contributions include:
@@
  • We demystify the serverless billing practice through these new characterization results and analyses, and discuss implications for designing future performant and cost-efficient serverless systems.
@@ 2. Serverless Billing Models and Practices
Pay-per-use is the common billing practice on serverless platforms. Billing models are the most direct determinant of the serverless cost as they convert billable resources (i.e., computing resources that are being billed for cloud users) into monetary charges that users immediately perceive. Billing models vary across platforms. In this section, we systematically deconstruct these billing practices to reveal how they shape the cost of serverless, reveal the underlying reasons for relevant billing practices, and discuss implications.
@@ 2.1. Overview of Serverless Billing Models
Table 1 summarizes the pay-per-use billing model on major serverless platforms listed in recent market reports (web, 2023b; MarketsandMarkets, 2024). While definitions of billable resources, wall-clock time, and pricing vary across different serverless platforms, most public serverless platforms bill a function invocation based on four factors: (1) billable wall-clock duration, (2) resource allocation amount and actual resource consumption over billable duration, (3) billing granularity and/or minimum billing cutoffs, and (4) a fixed fee associated with each invocation, which can be generally modeled as:
@@
where $T$ is the billable wall-clock time (e.g., wall-clock execution duration or turnaround time including initialization duration), $R_{ALLOC}$ is the set of billable computing resources that follow allocation-based billing (e.g., vCPUs, memory, GPU, and storage), $ALLOC(r)$ defines the allocation amount of billable resources $r$ over $T$, $R_{USG}$ is the set of billable resources to which consumption-based billing applies (e.g., network bandwidths and consumed CPU time of Cloudflare Workers), $USG(r)$ defines the absolute usage amount of billable resources $r$ over $T$, $G_{r}$ and $G_{T}$ define the billing granularity of resource $r$ and wall-clock time $T$ for rounding up or minimum billing cutoff (e.g., 128 MB and 100 ms), $C_{r}$ is the per-unit price of resource $r$, and $C_{0}$ is the fixed invocation fee.
Figure 1 illustrates the CPU and memory price on major serverless platforms presented in Table 1, which shows that the per-unit resource prices are often very similar across platforms. Following the serverless versus non-serverless cost comparison discussed in §1, this consistency in high per-unit resource prices indicates that the high price of serverless computing is not the result of any single provider’s billing strategy (AWS already offers some of the lowest per-unit resource prices).
Figure 1. The per-unit vCPUs and memory prices are generally similar across serverless platforms (as of 2025-05-15).
@@ 2.2. Coupled Control Knobs and Billable Resources
Control knobs and billable resources are tied closely, but allocated resources are always billed directly or indirectly: The billable resources in serverless billing models are usually mainly defined by the available control knobs. Public serverless platforms usually bill the computing resources they expose to users as control knobs. For example, AWS Lambda, Vercel Functions, and the Azure Functions Flexible Consumption allocate vCPUs in proportion to the allocated memory size. Some platforms, such as Huawei Function Compute, Azure Functions Consumption, and Oracle Functions, offer only a fixed set of memory sizes or fixed vCPU–memory pairs, rather than fine-grained configurations (e.g., per-MB memory configuration). In these cases, billing often appears to be based solely on memory allocation/usage, but the cost of CPU is embedded implicitly within the memory price. For instance, an AWS Lambda function with 1,769 MB of memory (which corresponds to 1 vCPU (web, 2025k)) incurs a charge of \$$2.8792\times 10^{-5}$ per second, while a GCP function provisioned with 1 vCPU and 1,769 MB of memory costs \$$2.8319\times 10^{-5}$ per second. The price per GB-second of memory on the platforms that mainly expose and bill memory control knobs usually closely matches what one would pay for memory and CPU on platforms that allow separate CPU allocations. Also, the ratio between the unit prices of CPU (in vCPU-seconds) and memory (in GB-seconds) on GCP, AWS Fargate (a container hosting platform that bills CPU and memory separately) (web, 2025u), and IBM Cloud Code Engine (function workloads) lies in a narrow range between 9 and 9.64, indicating a broad industry consensus on the relative value of vCPU versus memory.
On serverless platforms where vCPU and memory settings are relatively decoupled, CPU and memory are typically billed as two separate resources. However, even these platforms impose limits on how finely resources can be tuned. For example, Alibaba Cloud requires the ratio of vCPU to memory (in GB) to remain between 1:1 and 1:4, with step sizes of 0.05 vCPUs and 64 MB of memory (web, 2025q). Similarly, GCP imposes a minimum CPU allocation on the configured memory size (e.g., allocated memory of 512 MB must be configured with at least 0.333 vCPUs) (web, 2025l). These constraints on resource control knobs usually reflect an underlying function placement challenge: highly unbalanced CPU-to-memory combinations can fragment the resource capacity on host servers, potentially leading to higher deployment costs; e.g., through decreased deployment density (Cao et al., 2025; Li et al., 2023b), or higher scheduling delay waiting for placement (Joosen et al., 2025; Psychas and Ghaderi, 2018).
@@ 2.3. Inflation of Billable Resources
Figure 2. The billable resources can be multiple times higher than actual consumption on major serverless platforms.
@@
Billable resources are greatly inflated under wall-clock time allocation-based billing: Understanding the amount of billable resources under different billing models is critical to evaluating the cost of serverless platforms. To measure how much billable resources users pay on public serverless platforms, we analyze 66.1 million [^1] The first day of data in traces contains over 98 million requests. We exclude the requests reporting zero CPU usage. requests from the first day of data in Huawei serverless traces (web, 2024b; Joosen et al., 2025) and compute the billable vCPU time and the billable memory resources of each request under the billing models presented in Table 1. To avoid distortions from differences in per-unit prices on different platforms (however, they are mostly similar as discussed in §2.1), we report raw billable resources rather than cost in dollars. Figure 2 shows the distribution of these billable vCPU times and memories across requests under several representative billing models and resource allocation patterns, including proportional vCPU allocation (AWS Lambda), fixed vCPU-memory combinations (Huawei), wall-clock duration and resource usage rounding (GCP and Azure), and usage-based billing (Cloudflare Workers). As discussed in §2.2, CPU pricing is usually embedded for platforms with memory-only billing. Therefore, we include billable vCPU time for AWS.
@@
The gap between billed and actual resource usage quantifies the degree of inflated billable resources on current serverless platforms. Our analysis reveals that, under current models, billable vCPU time exceeds actual CPU usage by a factor of 1.02×1.02$\times$1.02 × (Cloudflare) up to 3.99×3.$9$ (GCP) on average, and billable memory exceeds real memory use by 1.95×1.95$\times$1.95 × (Azure) up to 5.49×5.49$\times$5.49 × (GCP) on average, among which usage-based billing (Cloudflare billable CPU and Azure billable memory) shows the lowest inflation. While differences in unit pricing shift the curve horizontally, these ratios remain the same (we do not compare absolute costs across platforms). Also, when mapping Huawei’s reported vCPU and memory allocations to AWS, we choose the larger of the two values to match its proportional vCPU allocation, which makes AWS billable resources slightly higher than Huawei.
Figure 3. Huawei serverless traces (web, 2024b; Joosen et al., 2025) show that serverless functions usually have low utilization of allocated resources. Billable resources are further inflated by inflexible resource control knobs as no clear linear relationships between CPU and memory utilization rates exist.
@@
One of the major driving factors of inflated billable resources is allocation-based wall-clock time billing. Even AWS with one of the finest billing granularities (i.e., 1 ms), charges billable vCPUs and memory, 2.62×2.62$\times$2.62 × and 3.67×3.67$\times$3.67 × higher than actual consumption on average. Functions rarely consume their full resource allocation (Joosen et al., 2023; Mahgoub et al., 2022), and wall-clock time includes periods when functions hold resources but remain idle or use little (e.g., blocking on remote API calls). Figure 3 illustrates resource usage relative to allocations. More than 42% of requests use less than 50% of the allotted CPU, and around 88% of requests use less than half of the allotted memory. The scatter plot of CPU and memory utilization shows a Pearson correlation of $0.397$, smaller than the value reported on serverless traces of Huawei private cloud (i.e., $0.6$) in 2023 (Joosen et al., 2023). The low correlation indicates that decoupling vCPU and memory configurations is crucial for reducing inflated costs. Inflexible allocations force developers to over-provision one resource to satisfy another bottleneck (Bilal et al., 2023).
@@
Although usage-based billing offers the lowest inflation in billable resources, it currently faces limitations and has not gained widespread adoption across providers. Cloudflare Workers is the only platform we studied that bills only on actual CPU time and aligns well with the actual CPU usage. However, it caps code artifact at 10 MB and memory at 128 MB. This is mainly designed for small, short, single-threaded JavaScript or WebAssembly (Wasm) tasks (under 1–2 ms) running on the V8 engine within their content delivery network (CDN) (web, 2025z; Alpers, 2025; web, 2025r), rather than general serverless workloads. Note that we do not argue that billing solely on the absolute amount of consumed resources is the only valid model in serverless computing. This is because resource allocation tied to wall-clock time, particularly for resources like memory (which is risky to overcommit (Tian et al., 2022)), significantly impacts function scheduling and deployment density. An ideal pay-per-use billing model is one that tracks real usage, exhibiting a perfect positive correlation statistically. There have been recent studies that aimed to move towards this ideal direction through dynamic or cooperative scheduling, or new billing models (Cao et al., 2025; Zhao et al., 2024; Pei et al., 2024; Liu and Niu, 2023).

### v2

# Demystifying Serverless Costs on Public Platforms: Bridging Billing, Architecture, and OS Scheduling

@@ Abstract.
Public cloud serverless platforms have attracted a large user base due to their high scalability, plug-and-play deployment model, and pay-per-use billing. However, compared to virtual machines and container hosting services, modern serverless offerings typically impose higher per-unit time and resource charges. Additionally, billing practices such as wall-clock time allocation-based billing, invocation fees, and usage rounding up can further increase costs.
This work, for the first time, holistically demystifies these costs by conducting an in-depth, top-down characterization and analysis from user-facing billing models, through request serving architectures, and down to operating system scheduling on major public serverless platforms. We quantify, for the first time, how current billing practices inflate billable resources up to $4.35\times$ beyond actual consumption. Also, our analysis reveals previously unreported cost drivers, such as operational patterns of serving architectures that create overheads, details of resource allocation during keep-alive periods, and OS scheduling granularity effects that directly impact both performance and billing. By tracing the sources of costs from billing models down to OS scheduling, we uncover the rationale behind today’s expensive serverless billing model and practices and provide insights for designing performant and cost-effective serverless systems.
Serverless Computing, Cloud Computing, Performance Measurements, Billing Models, OS Scheduling
submissionid: 992copyright: acmlicensedcopyright: ccccs: Computer systems organization Cloud computingccs: General and reference Measurementccs: Software and its engineering Scheduling
@@ 1. Introduction
Serverless computing has become one of the mainstream cloud computing paradigms, enabling developers to quickly deploy scalable and event-driven applications on the cloud without needing to manage the underlying infrastructure (Jonas et al., 2019; Datadog, 2023). Major cloud providers offer serverless computing solutions, such as AWS Lambda (Services, 2025g), Google Cloud (GCP) Run functions (Cloud, 2025g), Azure Functions (Azure, 2025g), IBM Cloud Code Engine (Cloud, 2025o), and Cloudflare Workers (Cloudflare, 2025a). Serverless computing stands out as the purest existing pay-per-use cloud model, offering automated scaling—from zero to thousands of instances in seconds—and fine-grained billing. As a result, it is often advertised as cost-efficient (Eismann et al., 2021b; Services, 2025f; Jonas et al., 2019; Wood, 2024; Cloudflare, 2025d).
@@
The widely acknowledged benefits of serverless architectures—such as high scalability, fine-grained pay-per-use billing, freedom from infrastructure management, and seamless integration with other cloud services—are not without associated costs (Eivy and Weinman, 2017; Hellerstein et al., 2018; Wen et al., 2023). In terms of the per-unit resource price, serverless offerings are often priced higher than other cloud computing paradigms, such as virtual machines (VMs) and containers running on container hosting platforms. We demonstrate this by comparing the price of AWS Lambda functions, AWS EC2 VMs, and AWS Fargate containers, all configured on identical ARM-based hardware in the us-east-2 region. We specifically chose ARM due to the diverse and performance-varying nature of AWS’s x86 processors, which complicates fair comparisons across services. An AWS Lambda function with 1 vCPU, 1,769 MB of memory, and 512 MB of ephemeral storage costs $\mathdollar 2.3034\times 10^{-5}$ per second (Services, 2025f), while a compute-optimized EC2 instance (c6g.medium) with 1 vCPU, 2 GB memory, and 1 GB storage and an AWS Fargate container with the identical resource allocation as EC2 cost only $\mathdollar 9.4753\times 10^{-6}$ and $\mathdollar 1.1003\times 10^{-5}$ per second, which are 41.1% and 47.8% of the AWS Lambda price. The cost of VMs can be further decreased by at least two times if using a burstable instance (e.g., AWS EC2 t4g.small flavor). Also, this comparison does not include the invocation fee of AWS Lambda, which is $\mathdollar 2\times 10^{-7}$ for each request, whereas EC2 instances and Fargate containers do not charge request fees. Additionally, our analysis of billing practices on major serverless platforms uncovers significant over-accounting (§2), showing that users can be charged for computing resources up to $4.35$ times greater than their actual usage.
@@
These observations motivate a fundamental research question: What makes serverless expensive? We argue that the root cause of the high unit prices and expensive billing practices in serverless lies in the architecture of modern serverless computing systems. Resource consumption and overhead incurred by the underlying runtimes and control plane for request serving, such as sandbox provisioning, isolation, request dispatch, and keep-alive, translate into higher per-unit charges passed on to serverless users. Additionally, some of our measurements of resource allocation patterns and performance behaviors on major serverless platforms point part of the execution costs and performance fluctuations to the underlying operating system (OS) scheduling mechanisms.
@@
Serverless Platform Billable Time Billable Resources Billing Granularity/Cutoffs Control Knobs and Steps
AWS Lambda (Services, 2025g, c, f)
@@
(Request-Based Billing) (Cloud, 2025g, l, f)
@@
Google Cloud Run
(Instance-Based Billing)∗∗∗ (Cloud, 2025g, l, f)
Wall-Clock
Instance Time
Allocated Memory and CPU 100 ms
Memory 1 MB
CPU 1 vCPU
@@ 1.5 GB memory and 1 vCPU)
Azure Functions Premium Plan∗∗∗
 (Azure, 2025g, h, e)
Wall-Clock
Instance Time
Allocated Memory and CPU
1 month
(minimum monthly cost applies)
CPU and Memory
(Fixed Combos)
@@
Huawei Cloud Function Graph (Cloud, 2025m)
@@
This table and related analysis in §2 focus on the most basic billable computing resources (i.e., CPU and memory). Other billable resources (e.g., storage, GPUs, and network bandwidths) may apply in practice. ∗∗AWS bills wall-clock turnaround time that includes initialization duration starting August 2025 (Gupta and Gebhart, 5 07). ∗∗∗Instance-based billing applies, where platforms charge for resource allocation over the function runtime instance lifespan regardless of requests.
Table 1. The billable models of major public serverless platforms. The notion of billable time, billable resources, and billing granularity varies across different serverless platforms (as of 2025-05-15).
To uncover these costs, a detailed analysis of current billing models together with measurements of the underlying serverless systems is required. Previous studies have characterized major serverless platforms in terms of architecture, performance, and resource management (Adzic and Chatley, 2017; Wang et al., 2018; Yu et al., 2020; Cvetković et al., 2023). However, as serverless computing evolves rapidly and more serverless offerings become available, some earlier measurements do not reflect or fully capture the latest billing scheme and operation patterns (e.g., serving architecture and keep-alive behaviors) of the public serverless computing platforms. In this work, we revisit some of the previous measurements and extend some of their performance and overhead characterization to fit modern serverless systems.
We adopt a top-down approach to analyze serverless costs. We start with user-facing billing models and conduct large-scale trace analysis on the billing scheme. Then, we analyze the performance, overhead, and resource allocation patterns of modern serverless request serving architectures. Finally, we investigate the impact of OS scheduling in detail. By tracing sources of costs from the billing model down to kernel scheduling, we provide the first comprehensive decomposition of serverless overhead and reveal the rationale behind current billing practices. For example, our large-scale, trace-based billing model analysis reveals significant bill inflation due to wall-clock allocation-based billing (§2.3), turnaround time billing (§2.4), rounding up of resource usage and execution duration, coarse billing granularity, and high invocation fees (Table 1 and §2.5). Also, we investigate the dual penalty of slowdowns and higher bills stemming from the multi-concurrency model (§3.1), high overheads of the HTTP-based request serving architecture (§3.2), and details of resource allocation during keep-alive (§3.3). Furthermore, we reveal the widespread CPU overallocation issue on public serverless platforms for the first time (§4). Specifically, our main contributions include:
@@
  • We demystify the serverless billing practice through these new characterization results and analyses, and discuss implications (labeled with I) for designing future performant and cost-efficient serverless systems.
We have made our artifact publicly available [^1] https://doi.org/10.5281/zenodo.17162822.
@@ 2. Serverless Billing Models and Practices
Pay-per-use is the common billing practice on serverless platforms. Billing models are the most direct determinants of the serverless cost as they convert billable resources (i.e., computing resources that are being billed for cloud users) into monetary charges that users immediately perceive. Billing models vary across platforms. In this section, we systematically deconstruct these billing practices to reveal how they shape the cost of serverless, reveal the underlying reasons for relevant billing practices, and discuss implications.
@@ 2.1. Overview of Serverless Billing Models
Table 1 summarizes the pay-per-use billing model on major serverless platforms listed in recent market reports (Datadog, 2023; MarketsandMarkets, 2024). While definitions of billable resources, wall-clock time, and pricing vary across different serverless platforms, most public serverless platforms bill a function invocation based on four factors: (1) billable wall-clock duration, (2) resource allocation amount and/or actual resource consumption over billable duration, (3) billing granularity and/or minimum billing cutoffs, and (4) a fixed fee associated with each invocation, which can be generally modeled as:
@@
where $T$ is the billable wall-clock time (e.g., wall-clock execution duration, turnaround time including initialization duration, or function instance lifespan), $R_{ALLOC}$ is the set of billable computing resources that follow allocation-based billing (e.g., vCPUs, memory, GPU, and storage), $ALLOC(r)$ defines the allocation amount of billable resources $r$ over $T$, $R_{USG}$ is the set of billable resources to which consumption-based billing applies (e.g., network bandwidths and consumed CPU time of Cloudflare Workers), $USG(r)$ defines the absolute usage amount of billable resources $r$ over $T$, $G_{r}$ and $G_{T}$ define the billing granularity of resource $r$ and wall-clock time $T$ for rounding up or minimum billing cutoff (e.g., 128 MB and 100 ms), $C_{r}$ is the per-unit price of resource $r$, and $C_{0}$ is the fixed invocation fee.
Depending on whether to use the whole function instance lifespan as the billable time, the billing model can generally be categorized into request-based billing (e.g., platforms other than the two listed with instance-based billing in Table 1) and instance-based billing (e.g., Azure Functions Premium Plan and Google Cloud Run with instance-based billing in Table 1). In request-based billing, each request is charged separately based on its execution duration (or turnaround time) and/or allocated/consumed resources during the billable period, while instance-based billing usually charges for provisioned resources on always-ready/scaled-out instances (resource allocation over instance lifespan) regardless of requests. On most platforms, users can enable instance-based billing by changing the billing setting or configuring provisioned concurrency, minimum instances, or scale-down delay for their functions (Cloud, 2025f; Services, 2025d). The fixed invocation fee ($C_{0}$) is usually not applied under instance-based billing.
Figure 1 illustrates the CPU and memory prices on major serverless platforms presented in Table 1, which shows that the per-unit resource prices are often very similar across platforms. Following the serverless versus non-serverless cost comparison discussed in §1, this consistency in high per-unit resource prices indicates that (I1) the high price of serverless computing is not the result of any single provider’s billing strategy (AWS already offers some of the lowest per-unit resource prices).
Figure 1. Resource (i.e., vCPU and memory) prices on major serverless platforms discussed in Table 1. The per-unit vCPUs and memory prices are generally similar across serverless platforms (as of 2025-05-15).
@@ 2.2. Coupled Control Knobs and Billable Resources
(I2) Control knobs and billable resources are tied closely, but allocated resources are always billed directly or indirectly: The billable resources in serverless billing models are usually mainly defined by the available control knobs. Public serverless platforms usually bill the computing resources they expose to users as control knobs. For example, AWS Lambda, Vercel Functions, and the Azure Functions Flexible Consumption allocate vCPUs in proportion to the allocated memory size. Some platforms, such as Huawei Function Compute, Azure Functions Consumption, and Oracle Functions, offer only a fixed set of memory sizes or fixed vCPU–memory pairs, rather than fine-grained configurations (e.g., per-MB memory configuration). In these cases, billing often appears to be based solely on memory allocation/usage, but the cost of CPU is embedded implicitly within the memory price. For instance, an AWS Lambda function with 1,769 MB of memory (which corresponds to 1 vCPU (Services, 2025c)) incurs a charge of $2.8792\times 10^{-5}$ per second, while a GCP function (first generation with request-based billing) provisioned with 1 vCPU and 1,769 MB of memory costs $2.8319\times 10^{-5}$ per second. The price per GB-second of memory on the platforms that mainly expose and bill memory control knobs usually closely matches what one would pay for memory and CPU on platforms that allow separate CPU allocations. Also, the ratio between the unit prices of CPU (in vCPU-seconds) and memory (in GB-seconds) on GCP, AWS Fargate (a container hosting platform that bills CPU and memory separately) (Services, 2025e), and IBM Cloud Code Engine (function workloads) lies in a narrow range between 9 and 9.64, indicating a broad industry consensus on the relative value of vCPU versus memory.
On serverless platforms where vCPU and memory settings are relatively decoupled, CPU and memory are typically billed as two separate resources. However, even these platforms impose limits on how finely resources can be tuned. For example, Alibaba Cloud requires the ratio of vCPU to memory (in GB) to remain between 1:1 and 1:4, with step sizes of 0.05 vCPUs and 64 MB of memory (Cloud, 2025c). Similarly, GCP imposes a minimum CPU allocation on the configured memory size (e.g., allocated memory of 512 MB must be configured with at least 0.333 vCPUs) (Cloud, 2025i). These constraints on resource control knobs usually reflect an underlying function placement challenge: highly unbalanced CPU-to-memory combinations can fragment the resource capacity on host servers, potentially leading to higher deployment costs; e.g., through decreased deployment density (Cao et al., 2025; Li et al., 2023b), or higher scheduling delay waiting for placement (Joosen et al., 2025; Psychas and Ghaderi, 2018).
@@ 2.3. Inflation of Billable Resources
(I3) Billable resources are greatly inflated under wall-clock time allocation-based billing: Understanding the amount of billable resources under different billing models is critical to evaluating the cost of serverless platforms. To measure how much billable resources users pay on public serverless platforms, we analyze 558.74 million [^2] The Huawei trace contains over 947.97 million requests. We exclude the requests reporting zero CPU usage or missing valid pod IDs/flavors. requests from the Huawei serverless trace (Huawei Public request tables) (Joosen, 2025; Joosen et al., 2025) and compute the billable vCPU time and the billable memory resources of each request under the billing models presented in Table 1. To avoid distortions from differences in per-unit prices on different platforms (however, they are mostly similar as discussed in §2.1), we report raw billable resources rather than cost in dollars. Figure 2 shows the distribution of these billable vCPU times and memories across requests under several representative billing models and resource allocation patterns, including proportional vCPU allocation (AWS Lambda), fixed vCPU-memory combinations (Huawei), wall-clock duration and resource usage rounding (GCP and Azure), and usage-based billing (Cloudflare Workers). As discussed in §2.2, CPU pricing is usually embedded for platforms with memory-based billing. Therefore, we include billable vCPU time for AWS.
Figure 2. Billable resources under different billing models. The billable resources can be multiple times higher than actual consumption on major serverless platforms.
@@
The gap between billed and actual resource usage quantifies the degree of inflated billable resources on current serverless platforms. Our analysis reveals that, under current models, billable vCPU time exceeds actual CPU usage by a factor of $1.01\times$ (Cloudflare) up to $3.63\times$ (GCP) on average, and billable memory exceeds real memory use by $1.57\times$ (Azure) up to $4.35\times$ (GCP) on average, among which usage-based billing (Cloudflare billable CPU and Azure billable memory) shows the lowest inflation. While differences in unit pricing shift the curve horizontally, these ratios remain the same (we do not compare absolute costs across platforms). Also, when mapping Huawei’s reported vCPU and memory allocations to AWS, we choose the larger of the two values to match its proportional vCPU allocation, which makes AWS billable resources slightly higher than Huawei.
@@
One of the major driving factors of inflated billable resources is allocation-based wall-clock time billing. Even AWS, with one of the finest billing granularities (i.e., 1 ms), charges billable vCPUs and memory, $2.49\times$ and $2.72\times$ higher than actual consumption on average. Functions rarely consume their full resource allocation (Joosen et al., 2023; Mahgoub et al., 2022b), and wall-clock time includes periods when functions hold resources but remain idle or use little (e.g., blocking on remote API calls). Figure 3 illustrates resource usage relative to allocations. More than 65% of requests use less than 50% of the allotted CPU, and around 76% of requests use less than half of the allotted memory. The scatter plot of CPU and memory utilization shows a Pearson correlation of $0.552$ and a Spearman correlation of 0.565, which is slightly smaller than the value reported on serverless traces of Huawei private cloud (i.e., $0.6$) in 2023 (Joosen et al., 2023). The moderate Pearson correlation suggests that the linear relationship is not strong, indicating that decoupling vCPU and memory configurations is crucial for reducing inflated costs. Inflexible allocations (e.g., proportional/linear) force developers to over-provision one resource to satisfy another bottleneck (Bilal et al., 2023; Verma et al., 2015).
Figure 3. Resource utilization rate distributions and their correlations. Huawei serverless traces (Joosen, 2025; Joosen et al., 2025) show that serverless functions usually have low utilization of allocated resources. Billable resources are further inflated by inflexible resource control knobs as no strong linear relationships between CPU and memory utilization rates exist.
@@
Although usage-based billing offers the lowest inflation in billable resources, it currently faces limitations and has not gained widespread adoption across providers. Cloudflare Workers is the only platform we studied that bills only on actual CPU time and aligns well with the actual CPU usage. However, it caps code artifacts at 10 MB and memory at 128 MB. This is mainly designed for small, short, single-threaded JavaScript or WebAssembly (Wasm) tasks (under 1–2 ms) running on the V8 engine within their content delivery network (CDN) (Cloudflare, 2025c; Alpers, 2025; Cloudflare, 2025b), rather than general serverless workloads. Note that we do not argue that billing solely on the absolute amount of consumed resources is the only valid model in serverless computing. This is because resource allocation tied to wall-clock time, particularly for resources like memory (which is risky to overcommit (Tian et al., 2022)), significantly impacts function scheduling and deployment density. An ideal pay-per-use billing model is one that tracks real usage, exhibiting a perfect positive correlation statistically. There have been recent studies that aimed to move towards this ideal direction through dynamic or cooperative scheduling, or new billing models (Cao et al., 2025; Zhao et al., 2024; Pei et al., 2024; Liu and Niu, 2023).